OpenStudy (anonymous):

Match the situation with the correct probability. For the marbles: a bag of 20 marbles has 8 blue, 2 red, 4 orange, and 6 purple. 1.P(rolling an odd number w/ a fair die 2.P(red marble) 3.P(drawing a red jack from a fair deck 4.P(marble) 5.P(rolling a 3 on a fair die 6.P(drawing an ace from a fair deck of cards 7.P(purple marble) 8.P(white marble)

OpenStudy (anonymous):

answers: A..5 B.1/13 C.1/26 D.1/6 E..1 F.0 G.1/4 H.1 I..3

OpenStudy (anonymous):

@arch1

OpenStudy (ganpat):

2. red = 2/20 = 1/10 3. 1/2 4. 20 / 20 = 1 :D 5. 1/6 6. 2/4 7. 6/20 = 3/10 8. 0.. as there is no white marble in the bag :D and whichever is missing.. is answer for 1st :-p

OpenStudy (ganpat):

hey, mistake in two options.. 3. 2/ 52 = 1/26.. 6. 4/52 = 1/13...

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

oh thanks i understand!

OpenStudy (anonymous):

@Ganpat you dnt have any of my answer choices i put the answers up there to choose from match them with the correct one

OpenStudy (anonymous):

@timo86m

OpenStudy (ganpat):

@arch1 : can u help..

OpenStudy (anonymous):

Thanks for tring though @Ganpat

OpenStudy (anonymous):

*trying

OpenStudy (ganpat):

i doubt ur options A and I i.e. 5 & 3 because as per my knowledge.. probability can never be more than 1.. because you cannot go select something more than what u have..

OpenStudy (anonymous):

der's no option lyk 1/10 for second que..bt according to me @ganpat is ryt

OpenStudy (anonymous):

its .5 not 5 and .3 not 3

OpenStudy (anonymous):

@arch1 and @Ganpat ^^^

OpenStudy (ganpat):

2. red = 2/20 = 1/10 = 0.1 -----> E 3. 1/26 --------------> C 4. 20 / 20 = 1 -------> H 5. 1/6 ------> D 6. 1/13 -----> B 7. 6/20 = 3/10 ------> I 8. 0.. as there is no white marble in the bag ---- > F not sure for 1st one... but as a guess.. 3/6 = 1/2 = 0.5 ----> A

OpenStudy (ganpat):

look man, you have all the options in my answer..

OpenStudy (ganpat):

@janessia : understood ??

OpenStudy (anonymous):

hold on let me take a look

OpenStudy (anonymous):

yes i understand! thanks so much!

OpenStudy (ganpat):

great.. so i believe next time, you gonna try and make it on ur own..

OpenStudy (anonymous):

of course!

OpenStudy (anonymous):

lol

OpenStudy (ganpat):

:)